OpenStudy (anonymous):

If f(x) = 5x + 3, then what is f-1(x) ?

OpenStudy (whpalmer4):

\[y = 5x+3\]Now swap x and y, and solve for y \[x = 5y+3\]\[y = f^{-1}(x) = ?\]

OpenStudy (anonymous):

no clue:((((((((

OpenStudy (whpalmer4):

Oh it's easy. \[x = 5y+3\]Solve that for y in terms of x, and the result is your \(f^{-1}(x)\). Hint, subtract 3 from both sides. Then...

OpenStudy (whpalmer4):

\[x - 3 = 5y + 3 - 3\]\[x - 3 = 5y\]\[y = ?\]